# {refsection}

### v1

@@
Fluid flow has a major effect on the aggregation and fragmentation of bacterial colonies. Yet, a generic framework to understand and predict how hydrodynamics affects colony size remains elusive. This study investigates how fluid flow affects the formation and maintenance of large colonial structures in cyanobacteria. We performed experiments on laboratory cultures and lake samples of the cyanobacterium Microcystis, while their colony size distribution was measured simultaneously by direct microscopic imaging. We demonstrate that EPS-embedded cells formed by cell division exhibit significant mechanical resistance to shear forces. However, at elevated hydrodynamic stress levels (exceeding those typically generated by surface wind mixing) these colonies experience fragmentation through an erosion process. We also show that single cells can aggregate into small colonies due to fluid flow. However, the structural integrity of these flow-induced colonies is weaker than that of colonies formed by cell division. We provide a mathematical analysis to support the experiments and demonstrate that a population model with two categories of colonies describes the measured size distributions. Our results shed light on the specific conditions wherein flow-induced fragmentation and aggregation of cyanobacteria are decisive and indicate that colony formation under natural conditions is mainly driven by cell division, although flow-induced aggregation could play a role in dense bloom events. These findings can be used to improve prediction models and mitigation strategies for toxic cyanobacterial blooms and also offer potential applications in other areas such as algal biotechnology or medical settings where the dynamics of biological aggregates play a significant role.
@@ Introduction
Toxic cyanobacterial blooms are a major nuisance for freshwater ecosystems worldwide [huisman2018cyanobacterial]. These blooms often lead to increased turbidity, cause harm to other aquatic organisms, and deteriorate water quality. Recent evidence indicates that the severity and frequency of cyanobacterial blooms are increasing due to eutrophication and climate change [paerl2009climate, visser2016rising, o2012rise, ho2019widespread]. Consequently, water management authorities show growing interest in techniques to prevent, monitor, and control these toxic blooms [ibelings2016cyanocost]. A particular effort is being taken to combat blooms of Microcystis, a non-motile colony-forming freshwater cyanobacterial genus that dominates many aquatic ecosystems worldwide [xiao2018colony, harke2016review]. Fluid dynamics is at the heart of Microcystis bloom formation, but also on some of the methods to control them. A large colony size and positive buoyancy give Microcystis a higher average light dose in comparison to non-buoyant algae [visser1997modelling, huisman2004changes]. Control methods such as deep artificial mixing can disrupt this competitive advantage by increasing the turbulence level in lakes [visser2016artificial]. Yet, the mechanisms of Microcystis colony formation, and in particular, the impact of fluid flow on the formation of colonies are poorly understood.
@@
Microcystis cells are held together by a layer of extracellular polymeric substances (EPS), facilitating the formation of large colonies of up to a millimeter in diameter with diverse morphological arrangements [van2022microcystis, liu2018characteristics, via2004distribution]. Two main mechanisms for colony formation have been proposed [xiao2018colony]: (i) cell division, i.e., clonal expansion of cells which remain attached by their encapsulating EPS layer after division, and (ii) aggregation, i.e., initially independent cells/colonies collide and adhere to each other. On the one hand, sequencing of the genetic diversity within Microcystis colonies supports the hypothesis that colony formation under natural conditions is primarily driven by cell division [perez2021single]. On the other hand, cell aggregation can occur on a shorter time scale and may offer improved protection against high grazing pressure [xiao2017meta]. Aggregation becomes significant when the collision frequency is high enough for aggregation rates to overcome cell division rate [burd2009particle], and the stickiness of colonies is sufficiently large (controlled by the EPS composition, pH, and ionic strength of water or medium [van1988aggregation, kiorboe1993phytoplankton, verspagen2006aggregation]). In addition to aggregation, fluid flow also influences colony formation through fragmentation, potentially serving as a mechanism for colony reproduction [day2022varied] to overcome the growth limitations of large colonies [feng2020structural]. Beyond modulating colony size through aggregation and fragmentation, fluid flow has also been proposed as a driving factor in changes in colony morphology and growth rate [li2018morphospecies, wilkinson2016effect, xiao2016effect].
@@
In the present study, we investigate to what extent Microcystis colony fragmentation and/or aggregation occurs at levels of turbulence observed for naturally or artificially mixed lakes. To this end, colonies of a laboratory culture of Microcystis were subjected to various intensities of shear flow (Fig. 1). We identified the mechanism of colony fragmentation using direct measurements of dynamic changes in colony size distribution under controlled laboratory conditions and analyze the characteristics of colonies formed by aggregation and cell division. Furthermore, the fragmentation behavior of the laboratory culture was compared with field samples of Microcystis spp. We also developed a numerical model based on a two-category discrete population to simulate changes in the size of Microcystis colonies via aggregation and fragmentation. Finally, we provide a phase map based on flow characteristics and cyanobacterial abundance to highlight different regimes in which fluid flow plays a significant role in cyanobacterial colony formation.
@@
A culture of Microcystis strain V163 was filtered to collect large colonies, and colonies on the filter were subsequently re-dispersed in fresh BG-11 medium up to the desired total biovolume fraction $\phi$ (Materials and Methods). The flow was generated by a cone-and-plate shear operating in an inertial regime, and it was characterized by the average energy dissipation rate $\dot{\varepsilon}$, hereafter called dissipation rate (Materials and Methods). The size distribution of the colonies was measured over time by microscopic imaging of the suspension through the transparent lower surface (Fig. 1). The colony size is described here by its relative diameter $l={d/d_{1}}$, where $d$ is the Feret’s diameter (maximum distance between two boundary points of the colony) and $d_{1}$ is the single-cell diameter. The number of cells in a colony, $N_{c}$, was calculated from its relative diameter via the relation $N_{c}=l^{f}$, where $f=2.09$ is the average fractal dimension measured for colonies of Microcystis strain V163 (Materials and Methods). The size distribution was computed as the normalized biovolume distribution $n(l)$, i.e., the number of cells in colonies in size bin $\{l;l+\Delta l\}$, divided by the total number of cells in the suspension.
@@
Direct measurement of the fragment distribution function $g(l,l^{\prime})$ (i.e., the frequency of fragments with size $l$ arising from the fragmentation of a colony of size $l^{\prime}$) is an experimental challenge [flesch1999laminar, vankova2007emulsification, haakansson2020experimental]. Individual fragmentation events could not be visualized as the residence time of colonies inside the field of view is too short. Instead, a suitable model can be identified from the rate of change in the biovolume distribution (Fig. 2E). A negative rate was observed for large colonies in the range $l\gtrsim 25$, indicating a loss in that size range by fragmentation. In contrast, the newly created fragments appeared as positive peaks, with a distinct peak close to the single-cell size. This result is compatible with an erosion mechanism, in which small fragments with one or a few cells are detached from the outer layer of the colony. For an ideal binary erosion, a fragmentation event of a colony with $i$ cells results in one cell being eroded from the surface of the colony with a complementary fragment of $i-1$ cells remaining. Small deviations from an ideal binary erosion are observed in Figure 2E, with most eroded fragments lying in the range $1\leq l\leq 2$. Nevertheless, ideal erosion provides a good approximation for the results. Accurate identification of complementary fragments from the rate of change in the biovolume distribution was more difficult, because the broad initial distribution of our suspension leaded to a broad distribution of complementary fragments. In this case, the rate of change in the biovolume distribution shifted from a net loss of eroded colonies at larger sizes ( $l\gtrsim 25$) to a net gain of complementary fragments at intermediate sizes ( $5\lesssim l\lesssim 20$).
@@
Assuming an ideal binary erosion mechanism, the fragmentation frequency $\kappa(l)$ was computed from the temporal dependence of the biovolume distribution (Materials and Methods). Here, $\kappa(l)$ is defined as the average number of fragmentation events (loss of a single cell) that one colony of size $l$ suffers per unit of time. The fragmentation frequency of Microcystis strain V163 was calculated as a function of colony size for three values of dissipation rate (Fig. 2F). The results show that the fragmentation frequency increases with colony size. Furthermore, an increase in dissipation rate sharply intensifies the erosion of colonies, with a two order of magnitude increase in $\kappa(l)$ when $\dot{\varepsilon}$ varies from $1.1$ to $9.6$ ${m^{2}/s^{3}}$.
@@
Fig. 2: Kinetics of fragmentation of Microcystis strain V163 colonies under cone-and-plate shear flow. The laboratory culture was filtered to select mainly large colonies, and the total biovolume fraction was adjusted to $\phi=10^{-4}$. Suspensions were subjected to an intense dissipation rate ($\dot{\varepsilon}=5.8~{}{m^{/}s^{3}}$) in panels A-E. (A) The initial size distribution of colonies, expressed as biovolume fraction of the relative colony diameter (normalized by single cell diameter). The size distribution had a bimodal shape, with large colonies (in yellow) and small colonies (in green, composed mostly of single cells, dimers, and some trimers, as depicted by the inset). (B) Median diameter of small and large colonies as a function of time. Bars indicate limits of 25th and 75th percentiles. (C) Most large colonies have been fragmented after 1 hour of shear flow, but the bimodal shape remained. (D) The size distribution shifted from large to small colonies as the shear flow induced fragmentation. (E) The rate of change in biovolume distribution at $t=0$h. Negative values indicate loss of colonies by fragmentation, while positive values indicate newly created fragments. The distribution suggests an erosion mechanism, as depicted by the cartoon inside the plot. (F) Fragmentation frequency as a function of the relative diameter of colonies for three values of dissipation rate. The inset shows details for the low dissipation rates. Error bars indicate the standard deviation. Dashed lines and shaded regions in panels B-D indicate predictions from the population model given by Eq. (1), and the lines in panel F indicate the predictions by Eq. (4). Best fit parameters: $\alpha_{1}=0.023$, $S_{1}=0.034$ , $q_{1}=4.5$, $S_{2}=31$ , $q_{2}=4.1$.
@@
In addition to inducing fragmentation, a turbulent flow may also increase the collision frequency of colonies. If the collisions are adhesive and the suspension is concentrated enough, then new colonies will be formed by aggregation. We subjected a single-cell suspension to the cone-and-plate shear flow to investigate whether aggregation can be a dominant colony-forming mechanism compared to cell division. Cultures of the Microcystis strain V163 were filtered to obtain a single-cell suspension, and the total biovolume fraction was adjusted with fresh BG-11 medium (Materials and Methods). A typical initial bio-volume fraction distribution is presented in Figure 3A. In addition to single cells, some small colonies up to $l=3$ were present in the filtered suspension due to imperfect filtering and/or aggregation during the pre-stirring. The size distribution was unimodal (single peak), in contrast with the bimodal distribution (double peak) in the suspension of division-formed colonies (Fig. 2A). After about 1 hour at a moderate dissipation rate ($\dot{\varepsilon}=0.019~{}{m^{2}/s^{3}}$) and a total biovolume fraction of $\phi=10^{-4}$, the single-cell suspension showed aggregation of cells into small colonies, with a steady-state median diameter of $l=1.8$ (Figs. 3B and 3C). Note that cell division can be neglected as the measurement time was much shorter than the division time. An increase in the collision rates due to a single-cell suspension led to an increase in the median diameter, although the increase was minor. At a high total biovolume fraction ($\phi=5\cdot 10^{-4}$), a steady-state median diameter of $l=2.4$ was reached, with a negligible fraction of colonies above $l\gtrsim 5$ (SI Appendix, Fig. S4C). Therefore, the aggregation of single cells is likely limited by weak bonds between cells. At the steady-state regime of median diameter, the newly-formed aggregates with a large diameter would be quickly fragmented back into smaller parts.
@@
The results above indicate that bonds formed during flow-induced aggregation (whether between two single cells or two division-formed colonies) are weak and cannot withstand intense dissipation rates. In contrast, the bonds between cells within a single division-formed colony are significantly more resistant. This difference in mechanical strength of the cellular binding in aggregation-formed versus division-formed colonies can be attributed to the structure of the EPS layer that surrounds the colonies. Aggregated colonies would have interstices between cells that are not filled with EPS. In contrast, cell division allows enough time for the gaps to be filled with secreted EPS, effectively increasing the bond strength between cells (Figs. 3B and 3D). However, we could not confirm this hypothesis visually, as the EPS layer in Microcystis strain V163 is too thin ($<1\mu m$ ) for a clear identification using optical microscopy.
Fig. 3: Kinetics of aggregation for a single-cell suspension of Microcystis strain V163 at a moderate dissipation rate of $\dot{\varepsilon}=0.019~{}{m^{2}/s^{3}}$ and a total biovolume fraction of $\phi=10^{-4}$. (A) Initial size distribution of the suspension as a function of the relative diameter, composed mostly of single cells. (B) Median diameter of colonies formed by aggregation of single cells as a function of time. Bars indicate limits of 25th and 75th percentiles. (C) After 1 hour of shear flow, the size distribution has shifted towards slightly larger diameters. (D) Time behavior of a suspension of large division-formed colonies under the same moderate dissipation rate and total biovolume fraction. The bimodal size distribution is separated into large (yellow) and small (green) colonies. Dashed lines and shaded regions in panels B-D indicate predictions from the population model given by Eq. (1). Best fit parameters: $\alpha_{1}=0.023$, $S_{1}=0.034$ , $q_{1}=4.5$, $S_{2}=31$ , $q_{2}=4.1$.
@@
Our experimental results for colonies of Microcystis strain V163 (Fig. 3) suggest the presence of two categories of colonies: aggregated colonies (denoted as $C_{1}$) and division-formed colonies ($C_{2}$). The total biovolume distribution $n(l)$ is given by the sum of both distributions $n_{1}(l)+n_{2}(l)$, normalized such that $\int_{0}^{\infty}n(l)\,\mathrm{d}l=1$. Within times shorter than the cell division time, the biovolume distribution of each colony category is described by a population model [van1988aggregation]:
@@
where $A_{i}(l,t)$ and $F_{i}(l,t)$ describe the aggregation and fragmentation rates as a function of colony size for each of the two categories, and $T_{i}(l,t)$ is a transfer rate of biovolume between the two categories. The turbulent shear-induced collisions between colonies shift the distribution towards larger aggregates, and the aggregation rate is given by
@@
where $l_{o}=(l^{f}-l^{\prime f})^{1/f}$. The collision frequency is described by the kernel $\beta(l,l^{\prime})$ and for a turbulent shear flow, we have $\beta(l,l^{\prime})=0.163(l+l^{\prime})^{3}\sqrt{\dot{\varepsilon}/\nu}$ , where $\nu$ is kinematic viscosity of the fluid [saffman1956collision]. The ratio of aggregative collisions over total collisions is given by the stickiness parameter $\alpha_{i}$. For the category $C_{1}$ (aggregated colonies), the stickiness is positive, $\alpha_{1}>0$. In contrast, we neglect the aggregation of $C_{2}$ colonies (collisions between $C_{2}$-$C_{2}$ and $C_{1}$-$C_{2}$), since the bonds formed by aggregation would be weaker and quickly rupture again, such that $\alpha_{2}=0$. This assumption is corroborated by the absence of aggregation between large colonies in our experiments (Fig. 3D).
The colony formation will be limited by fragmentation if the colony exceeds a critical size $l^{*}_{i}=({\dot{\varepsilon}/S_{i}})^{-q_{i}/f}$, where $S_{i}$ and $q_{i}$ are constitutive parameters that depend on the aggregate structure and bond strength, with a lower bound at $l^{*}_{i}\geq 1$ [baebler2008modelling, zaccone2009breakup]. We assume that aggregation-formed colonies have a small critical size, as the overlap between the EPS layer of each cell is narrow, leading to weak bonds. Meanwhile, division-formed colonies have an EPS layer that fills the pores between cells, thus increasing the resistance to hydrodynamic stress. The fragmentation rate is given by a first-order kinetic equation [baebler2008modelling]:
@@
For an initial suspension composed mainly of large colonies at an intense dissipation rate ($\dot{\varepsilon}=5.8~{}{m^{/}s^{3}}$), ), the model prediction of a decrease in the median size of $C_{2}$ colonies up to a steady-state diameter is in good agreement with the experimental findings (Fig. 2B). Moreover, the erosion mechanism accurately predicts the biovolume transfer from $C_{2}$ to $C_{1}$ colonies (Fig. 2D). The use of the erosion mechanism for the fragment distribution function is further corroborated by previous studies simulating bacterial aggregates [byrne2011postfragmentation]. For $C_{1}$ colonies at this intense dissipation rate, the model predicts that their median diameter remains near $l^{*}=1$ (i.e., the lower bound for the critical size). However, the experimental results show a slightly wider size distribution of both the small and large colonies than predicted by an ideal erosion mechanism (Fig. 2C). Nevertheless, the global behavior of the size distribution is well described by the two-category model. Furthermore, a very good agreement is also observed between the measured fragmentation frequency and the model predictions given by Eq.(4) (Fig. 2F).

### v2

@@
Fluid flow has a major effect on the aggregation and fragmentation of bacterial colonies. Yet, a generic framework to understand and predict how hydrodynamics affects colony size remains elusive. This study investigates how fluid flow affects the formation and maintenance of large colonial structures in cyanobacteria, using an experimental technique that precisely controls hydrodynamic conditions. We performed experiments on laboratory cultures and lake samples of the cyanobacterium Microcystis, while their colony size distribution was measured simultaneously by direct microscopic imaging. We demonstrate that EPS-embedded cells formed by cell division exhibit significant mechanical resistance to shear forces. However, at elevated hydrodynamic stress levels (exceeding those typically generated by surface wind mixing) these colonies experience fragmentation through an erosion process. We also show that single cells can aggregate into small colonies due to fluid flow. However, the structural integrity of these flow-induced colonies is weaker than that of colonies formed by cell division. We provide a mathematical analysis to support the experiments and demonstrate that a population model with two categories of colonies describes the measured size distributions. Our results shed light on the specific conditions wherein flow-induced fragmentation and aggregation of cyanobacteria are decisive and indicate that colony formation under natural conditions is mainly driven by cell division, although flow-induced aggregation could play a role in dense bloom events. These findings can be used to improve prediction models and mitigation strategies for toxic cyanobacterial blooms and also offer potential applications in other areas such as algal biotechnology or medical settings where the dynamics of biological aggregates play a significant role.
@@ Introduction
Toxic cyanobacterial blooms are a major nuisance for freshwater ecosystems worldwide [huisman2018cyanobacterial]. These blooms often lead to increased turbidity, cause harm to other aquatic organisms, and deteriorate water quality. Recent evidence indicates that the severity and frequency of cyanobacterial blooms are increasing due to eutrophication and climate change [paerl2009climate, visser2016rising, o2012rise, ho2019widespread]. Consequently, water management authorities show growing interest in techniques to prevent, monitor, and control these toxic blooms [ibelings2016cyanocost]. A particular effort is being taken to combat blooms of Microcystis, a non-motile colony-forming freshwater cyanobacterial genus that dominates many aquatic ecosystems worldwide [xiao2018colony, harke2016review]. Fluid dynamics is at the heart of Microcystis bloom formation, but is also used in some of the methods to control them. A large colony size and positive buoyancy give Microcystis a higher average light dose in comparison to non-buoyant algae [visser1997modelling, huisman2004changes]. Control methods such as deep artificial mixing have been successfully applied in several lakes to disrupt this competitive advantage by increasing the turbulence level [visser2016artificial]. Yet, the mechanisms of Microcystis colony formation, and in particular, the impact of fluid flow on the formation of colonies are poorly understood.
@@
Microcystis cells are held together by a layer of extracellular polymeric substances (EPS), facilitating the formation of large colonies of up to a millimeter in diameter with diverse morphological arrangements [van2022microcystis, liu2018characteristics, via2004distribution]. Two main mechanisms for colony formation have been proposed [xiao2018colony]: (i) cell division, i.e., clonal expansion of cells which remain attached by their secreted EPS layer after division, and (ii) aggregation, i.e., initially independent cells/colonies collide and adhere to each other. Sequencing of Microcystis colonies has revealed a much higher genetic diversity between independent colonies than within colonies, thus supporting the hypothesis that colony formation under natural conditions is primarily driven by cell division [perez2021single]. However, cell aggregation is a possible route for increase in colony size on a short time scale, and may offer improved protection against high grazing pressure [xiao2017meta]. Aggregation becomes significant when the collision frequency is high enough for aggregation rates to overcome cell division rate [burd2009particle], and the stickiness of colonies is sufficiently large (controlled by the EPS composition, pH, and ionic strength of water or medium [van1988aggregation, kiorboe1993phytoplankton, verspagen2006aggregation]). Yet, it remains an open question to what extent aggregation of Microcystis cells and colonies is a relevant process in natural systems. In addition to aggregation, fluid flow also influences colony formation through fragmentation, potentially serving as a mechanism for colony reproduction [day2022varied] to overcome the growth limitations of large colonies [feng2020structural]. Beyond modulating colony size through aggregation and fragmentation, fluid flow has also been proposed as a driving factor in changes in colony morphology and growth rate [li2018morphospecies, wilkinson2016effect, xiao2016effect].
@@
In the present study, we investigated to what extent Microcystis colony fragmentation and/or aggregation occurs at levels of turbulence observed for naturally or artificially mixed lakes. To this end, colonies of a laboratory culture of Microcystis were subjected to various intensities of shear flow (Fig. 1). We identified the mechanism of colony fragmentation using direct measurements of dynamic changes in colony size distribution under controlled laboratory conditions and analyze the characteristics of colonies formed by aggregation and cell division. Furthermore, the fragmentation behavior of the laboratory culture was compared with field samples of Microcystis spp. We also developed a numerical model based on a two-category discrete population to simulate changes in the size of Microcystis colonies via aggregation and fragmentation. Finally, we provide a phase map based on flow characteristics and cyanobacterial abundance to highlight different regimes in which fluid flow plays a significant role in cyanobacterial colony formation.
@@
A culture of Microcystis strain V163 was filtered to collect large colonies, and colonies on the filter were subsequently re-dispersed in fresh BG-11 medium up to the desired total biovolume fraction $\phi$ (Materials and Methods). The flow was generated by a cone-and-plate shear operating in an inertial regime, and it was characterized by the average energy dissipation rate $\dot{\varepsilon}$, hereafter called dissipation rate (Materials and Methods). The size distribution of the colonies was measured over time by microscopic imaging of the suspension through the transparent lower surface (Fig. 1). The colony size is described here by its relative diameter $l={d/d_{1}}$, where $d$ is the Feret’s diameter (maximum distance between two boundary points of the colony) and $d_{1}$ is the single-cell diameter. The number of cells in a colony, $N_{c}$, was calculated from its relative diameter via the relation $N_{c}=l^{f}$, where $f=2.09$ is the average fractal dimension measured for colonies of Microcystis strain V163 (Materials and Methods). The size distribution was computed as the normalized biovolume distribution $n(l)$, i.e., the number of cells in colonies in size bin $\{l;l+\Delta l\}$, divided by the total number of cells in the suspension and the bin width.
@@
Direct measurement of the fragment distribution function $g(l,l^{\prime})$ (i.e., the frequency of fragments with size $l$ arising from the fragmentation of a colony of size $l^{\prime}$) is an experimental challenge [flesch1999laminar, vankova2007emulsification, haakansson2020experimental]. Individual fragmentation events could not be visualized as the residence time of colonies inside the field of view is too short. Instead, a suitable model can be identified from the rate of change in the biovolume distribution (Fig. 2E). A negative rate was observed for large colonies in the range $l\gtrsim 25$, indicating a loss in that size range by fragmentation. In contrast, the newly created fragments appeared as positive peaks, with a distinct peak close to the single-cell size. This result is compatible with an erosion mechanism, in which small fragments with one or a few cells are detached from the outer layer of the colony. For an ideal binary erosion, a fragmentation event of a colony with $i$ cells results in one cell being eroded from the surface of the colony with a complementary fragment of $i-1$ cells remaining. Small deviations from an ideal binary erosion are observed in Figure 2E, with most eroded fragments lying in the range $1\leq l\leq 2$. Nevertheless, ideal erosion provides a good approximation for the results. Accurate identification of complementary fragments from the rate of change in the biovolume distribution was more difficult, because the broad initial distribution of our suspension led to a broad distribution of complementary fragments. In this case, the rate of change in the biovolume distribution shifted from a net loss of eroded colonies at larger sizes ( $l\gtrsim 25$) to a net gain of complementary fragments at intermediate sizes ( $5\lesssim l\lesssim 20$).
@@
Assuming an ideal binary erosion mechanism, the fragmentation frequency $\kappa(l)$ was computed from the temporal dependence of the biovolume distribution (Materials and Methods). Here, $\kappa(l)$ is defined as the average number of fragmentation events (loss of a single cell) that one colony of size $l$ suffers per unit of time. The fragmentation frequency of Microcystis strain V163 was calculated as a function of colony size for three values of dissipation rate (Fig. 2F). The results show that the fragmentation frequency increases with colony size. Furthermore, an increase in dissipation rate sharply intensifies the erosion of colonies, with a two order of magnitude increase in $\kappa(l)$ when $\dot{\varepsilon}$ varies from $1.1$ to $9.6$ ${m^{2}/s^{3}}$. The intensified fragmentation frequency also reduces the time required for the biovolume distribution to reach a steady-state (SI Appendix, Fig. S3). It is also noticeable that some of the large colonies (5 to 10 % of the total biovolume) do not fragment to single cells but remain as large colonies of 10-20 cells even after 5 hours of very high dissipation rate ($\dot{\varepsilon}=9.6~{}{m^{2}/s^{3}}$), indicating variability in the mechanical strength of colonies, with a subset presenting an exceptional resistance in comparison to the average behavior (SI Appendix, Fig. S3E and F)
@@
Fig. 2: Kinetics of fragmentation of Microcystis strain V163 colonies under cone-and-plate shear flow. The laboratory culture was filtered to select mainly large colonies, and the total biovolume fraction was adjusted to $\phi=10^{-4}$. Suspensions were subjected to an intense dissipation rate ($\dot{\varepsilon}=5.8~{}{m^{2}/s^{3}}$) in panels A-E. (A) The initial size distribution of colonies, expressed as biovolume fraction of the relative colony diameter (normalized by single cell diameter). The size distribution had a bimodal shape, with large colonies (in yellow) and small colonies (in green, composed mostly of single cells, dimers, and some trimers, as depicted by the inset). (B) Median diameter of small and large colonies as a function of time. Symbols indicate the experimental data, while the lines indicate the predictions from the population model given by Eq. (1) with an erosion (dashed) or equal fragments (dot-dash) hypothesis for large division-formed colonies. Small aggregates colonies follow the equal fragments hypothesis. Bars and shaded region indicate limits of 25th and 75th percentiles. (C) Most large colonies have been fragmented after 1 hour of shear flow, but the bimodal shape of the colony size distribution remained. (D) Biovolume fraction of small colonies (i.e., biovolume of small colonies over the total biovolume) as a function of time. A shift is observed from large to small colonies, captured well by the erosion model, but not by the equal fragments model. (E) The rate of change in biovolume distribution at $t=0$h. Negative values indicate loss of colonies by fragmentation, while positive values indicate newly created fragments. The distribution suggests an erosion mechanism, as depicted by the cartoon inside the plot. (F) Fragmentation frequency as a function of the relative diameter of colonies for three values of dissipation rate. The inset shows details for moderate dissipation rates. Error bars indicate the standard deviation and the lines indicate the predictions by Eq. (5). Best fit parameters: $\alpha_{1}=0.023$, $S_{1}=0.034$ , $q_{1}=4.5$; erosion: $S_{2}=31$ , $q_{2}=4.1$ ; equal fragments: $S_{2}=33$ , $q_{2}=6.3$.
@@
In addition to inducing fragmentation, a turbulent flow may also increase the collision frequency of colonies. If the collisions are adhesive and the suspension is concentrated enough, then new colonies will be formed by aggregation. We subjected a single-cell suspension to the cone-and-plate shear flow to investigate whether aggregation can be a dominant colony-forming mechanism compared to cell division. Cultures of the Microcystis strain V163 were filtered to obtain a single-cell suspension, and the total biovolume fraction was adjusted with fresh BG-11 medium (Materials and Methods). A typical initial bio-volume fraction distribution is presented in Figure 3A. In addition to single cells, some small colonies up to $l=3$ were present in the filtered suspension due to imperfect filtering and/or aggregation during the pre-stirring. The size distribution was unimodal (single peak, Fig. 3A), in contrast with the bimodal distribution (double peak, Fig. 2A) in the suspension of division-formed colonies . After about 1 hour at a moderate dissipation rate ($\dot{\varepsilon}=0.019~{}{m^{2}/s^{3}}$) and a total biovolume fraction of $\phi=10^{-4}$, the single-cell suspension showed aggregation of cells into small colonies, with a steady-state median diameter of $l=1.8$ (Figs. 3B and 3C). Note that cell division can be neglected as the measurement time was much shorter than the division time. An increase in the collision rates due to a denser single-cell suspension led to an increase in the median diameter, although the increase was minor. At a high total biovolume fraction ($\phi=5\cdot 10^{-4}$), a steady-state median diameter of $l=2.4$ was reached, with a negligible fraction of colonies above $l\gtrsim 5$ (SI Appendix, Fig. S4C). Therefore, the aggregation of single cells is likely limited by weak bonds between cells. At the steady-state regime of median diameter, the newly-formed aggregates with a large diameter would be quickly fragmented back into smaller parts.
@@
The results above indicate that bonds formed during flow-induced aggregation (whether between two single cells or two division-formed colonies) are weak and cannot withstand intense dissipation rates. In contrast, the bonds between cells within a single division-formed colony are significantly more resistant. This difference in mechanical strength of the cellular binding in aggregation-formed versus division-formed colonies can be attributed to the structure of the EPS layer that surrounds the colonies. Aggregated colonies would have intercellular gaps (smaller that the cell size) that are not filled with EPS. In contrast, cell division allows enough time for the gaps to be filled with secreted EPS, effectively increasing the bond strength between cells (Figs. 3B and 3D). However, we could not confirm this hypothesis visually, as the EPS layer in Microcystis strain V163 is too thin ($<1\mu m$ ) for a clear identification using optical microscopy.
Fig. 3: Kinetics of aggregation for a single-cell suspension of Microcystis strain V163 at a moderate dissipation rate of $\dot{\varepsilon}=0.019~{}{m^{2}/s^{3}}$ and a total biovolume fraction of $\phi=10^{-4}$. (A) Initial size distribution of the suspension as a function of the relative diameter, composed mostly of single cells. (B) Median diameter of colonies formed by aggregation of single cells as a function of time. Symbols indicate the experimental data, while the lines indicate the predictions from the population model given by Eq. (1) with an equal fragment hypothesis for small aggregates colonies. Bars and shaded region indicate limits of 25th and 75th percentiles. (C) After 1 hour of shear flow, the size distribution has shifted towards slightly larger diameters. (D) Time behavior of a suspension of large division-formed colonies under the same moderate dissipation rate and total biovolume fraction. The bimodal size distribution is separated into large (yellow) and small (green) colonies. Dashed lines and shaded regions indicate predictions from the population model given by Eq. (1). Best fit parameters: $\alpha_{1}=0.023$, $S_{1}=0.034$ , $q_{1}=4.5$, $S_{2}=31$ , $q_{2}=4.1$.
@@
Our experimental results for colonies of Microcystis strain V163 (Fig. 3) suggest the presence of two categories of colonies: single cells and small aggregates (denoted as $C_{1}$) and large colonies formed by cell division ($C_{2}$). The total biovolume distribution $n(l)$ is given by the sum of both distributions $n_{1}(l)+n_{2}(l)$, normalized such that $\int_{0}^{\infty}n(l)\,\mathrm{d}l=1$. Within times shorter than the cell division time, the biovolume distribution of each colony category is described by a population model [van1988aggregation]:
@@
where $A_{i}(l,t)$ and $F_{i}(l,t)$ describe the aggregation and fragmentation rates as a function of colony size for each of the two categories ($i\in[1,2]$), and $T_{i}(l,t)$ is a transfer rate of biovolume between the two categories. The turbulent shear-induced collisions between colonies shift the distribution towards larger aggregates, and the aggregation rate is given by
@@
where $l_{o}=(l^{f}-l^{\prime f})^{1/f}$. The collision frequency is described by the kernel $\beta(l,l^{\prime})$ and for a turbulent shear flow, we have $\beta(l,l^{\prime})=0.163(l+l^{\prime})^{3}\sqrt{\dot{\varepsilon}/\nu}$ , where $\nu$ is kinematic viscosity of the fluid [saffman1956collision]. The ratio of aggregative collisions over total collisions is given by the stickiness parameter $\alpha_{i}$. For the category $C_{1}$ (single cells and small aggregates), the stickiness is positive, $\alpha_{1}>0$. In contrast, we neglect the aggregation of $C_{2}$ colonies (collisions between $C_{2}$-$C_{2}$ and $C_{1}$-$C_{2}$), since the bonds formed by aggregation would be weaker and quickly rupture again, such that $\alpha_{2}=0$. This assumption is corroborated by the absence of aggregation between large colonies in our experiments (Fig. 3D).
The colony formation will be limited by fragmentation if the colony exceeds a critical size $l^{*}_{i}$, given by
$l^{*}_{i}=\left(\dfrac{\dot{\varepsilon}}{S_{i}}\right)^{-q_{i}/f},$ (3)
where $S_{i}$ and $q_{i}$ are constitutive parameters that depend on the aggregate structure and bond strength, with a lower bound at $l^{*}_{i}\geq 1$ [baebler2008modelling, zaccone2009breakup]. We assume that aggregation-formed colonies have a small critical size, as the overlap between the EPS layer of each cell is narrow, leading to weak bonds. Meanwhile, division-formed colonies produce sufficient EPS to fill the space between cells, thus increasing the resistance to hydrodynamic stress. The fragmentation rate is given by a first-order kinetic equation [baebler2008modelling]:
@@
For an initial suspension composed mainly of large colonies at an intense dissipation rate ($\dot{\varepsilon}=5.8~{}{m^{2}/s^{3}}$), the model prediction of a decrease in the median size of $C_{2}$ colonies up to a steady-state diameter is in good agreement with the experimental findings (Fig. 2B). Moreover, the erosion mechanism accurately predicts the biovolume transfer from large to small colonies (Fig. 2D). In comparison, the hypothesis of equal fragments formation for $C_{2}$ colonies does not capture such transfer, as all fragments remain in the large colony range (Fig. 2D and Fig. S9 of SI Appendix).The use of the erosion mechanism for the fragment distribution function is further corroborated by previous studies simulating bacterial aggregates [byrne2011postfragmentation]. For $C_{1}$ colonies at this intense dissipation rate, the model predicts that their median diameter remains near $l^{*}=1$ (i.e., single cells). However, the experimental results show a slightly wider size distribution of both the small and large colonies than predicted by an ideal erosion mechanism (Fig. 2C). Nevertheless, the global behavior of the size distribution is well described by the two-category model. Furthermore, a very good agreement is also observed between the measured fragmentation frequency and the model predictions given by Eq. (5) (Fig. 2F).